OpenStudy (anonymous):

Using complete sentences, explain how to find the equation of the line, in standard form and slope-intercept form, passing through (-1, -7) and (1, -1). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope-intercept form. Will give medals!

OpenStudy (anonymous):

slope= \[\frac{ y2-y1 }{ x2-x1 } ,\frac{ -7--1 }{ -1-1 }\] slope = 3 then (y -y0)=slope*(x-x0) y--7=3 ( x--1) or y--1=3 (x-1)